OpenStudy (anonymous):

What is the equation of the line that passes through (12, 4) and is perpendicular to the graph of y = -3 over 4x - 2?

OpenStudy (anonymous):

Giving medals also, and fan!

OpenStudy (jdoe0001):

what is the slope of \(\bf y = -\cfrac{3}{4}x-2\quad ?\)

OpenStudy (anonymous):

m=-3/4

OpenStudy (jdoe0001):

right, so the slope of a line PERPENDICULAR to that one will be the NEGATIVE RECIPROCAL of that, what does that mean? well NEGATIVE of -3/4 ? 3/4 RECIPROCAL of that? 4/3 so \((\bf \begin{array}{lllll} &x_1&y_1\\ &(16\quad ,&4) \end{array} \\\quad \\ slope = m= \cfrac{4}{3} \\ \quad \\ y-y_1=m(x-x_1)\quad \textit{plug in your values and solve for "y"} \)

OpenStudy (jdoe0001):

16? darn... typo

OpenStudy (jdoe0001):

\(\bf \begin{array}{lllll} &x_1&y_1\\ &(12\quad ,&4) \end{array} \\\quad \\ slope = m= \cfrac{4}{3} \\ \quad \\ y-y_1=m(x-x_1)\quad \textit{plug in your values and solve for "y"}\)

OpenStudy (anonymous):

y = -3/4x + 13?

OpenStudy (anonymous):

wait wait wait

OpenStudy (anonymous):

y = 4 over 3x - 12?

OpenStudy (jdoe0001):

\(\bf \begin{array}{lllll} &x_1&y_1\\ &(12\quad ,&4) \end{array} \\\quad \\ slope = m= \cfrac{4}{3} \\ \quad \\ y-y_1=m(x-x_1)\implies y-12=\cfrac{4}{3}(x-12)\\ \quad \\ y-12=\cfrac{4}{3}x-\cfrac{4\cdot 12}{3}\implies y-12=\cfrac{4}{3}x-\cfrac{48}{3}\implies y-12=\cfrac{4}{3}x-16\\ \quad \\ y=\cfrac{4}{3}x-16+12 \)

OpenStudy (jdoe0001):

\(\bf y-y_1=m(x-x_1)\implies y-12=\cfrac{4}{3}(x-12)\\ \quad \\ y-12=\cfrac{4}{3}x-\cfrac{4\cdot 12}{3}\implies y-12=\cfrac{4}{3}x-\cfrac{48}{3} \\ \quad \\y-12=\cfrac{4}{3}x-16\\ \quad \\ y=\cfrac{4}{3}x-16+12 \) anyhow, since it got a bit truncated

OpenStudy (anonymous):

So I was correct?

OpenStudy (jdoe0001):

well... .-16 + 12 \(\ne -12\)

OpenStudy (anonymous):

-12 is part of my answer :D